OpenStudy (anonymous):

45 is what % of 360?

OpenStudy (angela210793):

x/100 *360=45 36x=450 x=450/36=12.5

OpenStudy (anonymous):

12.5

OpenStudy (anonymous):

thanx

OpenStudy (anonymous):

338 is what % of 250?

OpenStudy (angela210793):

solve it in the same way Roseanne...try at lesat :)

OpenStudy (anonymous):

im confused to be honest how to solve it

OpenStudy (anonymous):

135.2

OpenStudy (anonymous):

What is 0.5 % of 150?

OpenStudy (angela210793):

0.5/100*150=7.5

OpenStudy (anonymous):

.75